OpenStudy (anonymous):

simplify : 2x^2+8x-42 over x^2+12x+35 he;lp please

OpenStudy (chaise):

\[\frac{2x^2+8x-42}{x^2+12x+35}\] \[=\frac{2(x-3)(x+7)}{(x+5)(x+7)}\] You can now cancel like terms. You get: =2(x-3)/x+5 =2x-6/x+5

OpenStudy (anonymous):

thank you